OpenStudy (anonymous):

Solve by using the perfect squares method. x2 + 4x + 4 = 0 2 –2 4 –4

OpenStudy (anonymous):

\(\Large x^2 + 4x+4=0 \Rightarrow x=\frac{-4\pm \sqrt{4^2-4\times4}}{2} = \frac{-4}{2}=-2\)